OpenStudy (anonymous):

Prove by contradiction. Suppose that x and y are positive integers. Show that the sq root of (x^2 + y^2) is not equal to x + y.

OpenStudy (anonymous):

proof by contraction means suppose \[\sqrt{x^2+y^2}=x+y\]

OpenStudy (anonymous):

then arrive at a contradiction, square both sides get \[x^2+y^2=(x+y)^2\]

OpenStudy (anonymous):

so \[x^2+y^2=x^2+y^2+2xy\] \[0=2xy\] imples either \(x=0\) or \(y=0\)

OpenStudy (anonymous):

and since 0 is not a positive integer, this contracts the assumption that both x and y are positive integers